OpenStudy (anonymous):

CHECK ANSWER:20 Use the binomial coefficient formula to calculate the binomial coefficient for the 4th term in a polynomial expansion when the power of the binomial is 6.

OpenStudy (anonymous):

are u sure because i got 20 and that is a choice on here and 15 is not

OpenStudy (anonymous):

sorry its 20

OpenStudy (anonymous):

its : \[\left(\begin{matrix}6 \\ 3\end{matrix}\right) x^{6-3}y^{3}= 20x^3y^3\]

OpenStudy (anonymous):

If i confused you am really sorry, I totally forgot that \[\left(\begin{matrix}6 \\ 0\end{matrix}\right)\] was the first term. instead I took \[\left(\begin{matrix}6 \\ 1\end{matrix}\right)\] as the first term. If there are any other ones i got wrong let me know so i can delete the post and put the correct one up